EOQ

Bundaberg Glass Company is a distributor of car windscreens. The windscreens are manufactured in Japan and shipped to Bundaberg. Management is expecting an annual demand of 10 800 windscreens. The purchase price of each windscreen is $400. Other costs associated with ordering and maintaining an inventory of these windscreens are shown below.

 

§  The ordering costs incurred in the purchase order department for placing and processing orders for the past three years are shown below.

 

Year                Orders placed and processed              Total processing costs

1                                  17                                                        $12300

2                                  55                                                        12475

3                                  97                                                        12700

 

§  Management expects these ordering costs to increase by 16 per cent over the amounts and rates experienced in the last three years.

§  Each order is inspected by Australian custom officers. A fee of $47 is charged.

§  A clerk in the receiving department receives, inspects and secures the windscreens as they arrive from the manufacturer. The activity requires 4 hours per order received. This clerk has no other responsibilities and is paid at the rate of $24 per hour. Related variable overhead costs in this department are applied at the rate of $6 per hour.

§  Additional warehouse space will have to be rented to store the new windscreens. Space can be rented as needed in a warehouse at an estimated cost of $2500 per year plus $6.25 per windscreen.

§  Breakage cost is estimated to average $4 per windscreen.

§  Insurance on the inventory costs $1.25 per windscreen.

§  Other carrying costs amount to $8.50 per windscreen.

 

Bundaberg Glass Company works 6-day week for 50 weeks each year. The firm is closed for two weeks each year. Six working days are required for the time the order is placed with the supplier until it is received.

 

Required:

 

1.     Assuming that all costs other than the order costs remain the same, calculate the following amounts for the Bundaberg Glass Company for year 4:

a.      The amount of the ordering cost that should be used in the EOQ formula (Hint: Use the high-low method to estimate the incremental processing cost per order.)

b.     Amount of the carrying cost that should be used in the EOQ formula

c.      Economic order quantity

d.     Minimum annual relevant cost of ordering and carrying at the economic order quantity

e.      Reorder point in units.

 

2.     Management has been able to negotiate a JIT purchasing agreement with the Japanese manufacturer, and the inspection fee has been renegotiated with the custom officials. The purchasing manager has determined that JIT purchasing would enable the company to reduce the ordering cost to $32.50 per order. Moreover, she has analysed the cost of storing windscreens, taking care to include the cost of wasted space and inefficiency. She estimates that the real annual cost of carrying inventory is $60 per windscreen.

a.      Calculate the new EOQ, given the purchasing manager’s new cost estimates

b.     How many orders would now be placed each year?

c.      Calculate the new minimum annual relevant cost of ordering and carrying inventory.
Answer #1

Ans. (a)

Calculation of ordering cost per order

Particulars$
order processing costs per order(Note-1)14,175.20
order inspection cost47.00
receiving department cost{$(24+6)×4 hours}120.00
Total ordering cost per order14,342.20

Note-1

order processing costs

Variable cost per order = $(12,700-12,300)/(97-17)orders

= $5 per order

Fixed cost = $12,700-(97×5)=$12,215

Total order processing costs per order = $5+$12,215=$12,220

Increase in order processing costs = 16%

Increased order processing costs per order = $12220×1.16=$14,175.20

Ans.(b)

Calculation of carrying cost per unit per annum

Particulars$
warehouse rent {($2,500/70,000)+$6.25}6.48
Breakage cost4.00
insurance costs1.25
other carrying costs8.50
Total carrying cost per unit per annum20.23

Ans.(c)

EOQ = (2AO/C)^1/2

where,

A= annual demand(10,800 units)

O=ordering cost per order ($14,342.20)

C=carrying cost per unit p.a.($20.23)

Accordingly,

EOQ={(2×10800×14342.20)/20.23}^1/2

EOQ=3913 units

Ans.(d)

At EOQ

(1)Calculation of minimum annual relevant ordering cost

Particulars$
order processing costs (Note-2)14,185.41
order inspection cost{(10800/3913)×47}129.72
receiving department cost{(10800/3913)×120}331.20
minimum annual relevant ordering cost14646.33

Note-2

Order processing cost

Fixed = $12,215×1.16=$14,169.40

variable cost per order = $5×1.16=$5.80

Variable cost = {(10800/3913)×$5.80}=$16.01

Total order processing cost=$14,169.40+$16.01=$14,185.41

(2)Calculation of minimum annual relevant carrying cost

=(3913/2)×$20.23

=$39580

Ans. (e)

Reorder point =Maximum Daily Usage x Maximum Lead Time Days

=36 units × 6 days

=216 units

Note-3

Maximum Daily Usage = 10800 units/(6×50)days

=36 units


answered by: play.gg